Title: Skew products which are not standard

Abstract: For any function f on the unit circle and T_{t} a flow, we can form the skew product extension T̂_{f}(x, ω) = (2x mod 1, T_{f(x)}(ω)). Like the T, T^{−1} map, this is a random walk on a random scenery. The talk will focus on the proof that for f satisfying a Hölder condition with ∫ f = 0 and T_{t} a positiveentropy flow, the sequence of random variables arising from T̂_{f} is not standard. The proof follows the method used by Heicklen and Hoffman to prove that the T, T^{−1} endomorphism is not standard.